[figure: 9.7 cm]

Analytical determination of resistivity components from multiterminal measurements in highly anisotropic media

P. Popovich^{a}, S. Mészáros^{b}, K. Vad^{b} and J. Hakl^{b} aUzhgorod National University, Voloshin str., 54, UA-88000 Uzhgorod, Ukraine bInstitute of Nuclear Research, PO. Box 51, H-4001 Debrecen, Hungary

We present a method to determine the resistivity tensor of materials with high uniaxial resistivity anisotropy (ρ_{c} ≫ ρ_{ab}). The method is based on analytical solution of the Laplace’s equation for parallelepiped shaped samples. The current contacts are located on opposite sides of sample and the input data we use are the voltage differencies on the top and bottom surfaces. The analytical equations for ρ_{ab} and ρ_{c} are given for high values of effective anisotropy Γ. The influence of the position and finite size of current contacts are taken into account. The advantages of our analytical solution are that it is faster than numerical simulations and the applied contact arrangement increases the accuracy of the measurement.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Paul Popovich

Address for correspondence : Physikalisches Institut - Experimentalphysik II, Universität Tübingen, Auf der Morgenstelle 14, D-72076 Tübingen, Germany

Email address : popovich@brahms.pit.physik.uni-tuebingen.de